OpenStudy (anonymous):

2^xsquare=4^2x-2 How do I solve this question,please help.

OpenStudy (solomonzelman):

Can you retype the question, i don't get it.

OpenStudy (solomonzelman):

(Use equation editor on the bottom left corner)

OpenStudy (anonymous):

\[2^{x2}=4^{2x-2}\]

OpenStudy (solomonzelman):

\[2x~Log~2=(2x-2)~Log~4\]\[2x/(2x-2)=Log~4/Log~2\]

OpenStudy (solomonzelman):

good till now?

OpenStudy (anonymous):

Yup,but is it possible to use the indices law to solve this?

OpenStudy (solomonzelman):

Yes yes

OpenStudy (solomonzelman):

\[Log_35=\frac{Log_{10}5}{Log_{10}3}\]see this example?

OpenStudy (anonymous):

Yes,I am fairly new to logarithm.Could you help me on that too?

OpenStudy (solomonzelman):

Do you see this example above get it? yes/no

OpenStudy (solomonzelman):

(it a formula/identity)

OpenStudy (anonymous):

I don't get it

OpenStudy (solomonzelman):

You are saying that it's = log of number over log of base.

OpenStudy (anonymous):

Is this the change of base law?

OpenStudy (solomonzelman):

YES.

OpenStudy (solomonzelman):

Good so far?

OpenStudy (anonymous):

yup

OpenStudy (solomonzelman):

look at the left right of you question.\[\frac{Log4}{Log2}\](presumably, when not spesified the base is 10) just like if you see \[\sqrt{?}~~~~~i t's ~~~~~\sqrt[2]{?}\], right?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

So it would be the same thing as \[\frac{Log_{10}4}{Log_{10}2}\]

OpenStudy (solomonzelman):

\[\frac{Log_{10}4}{Log_{10}2}=Log_24\]So far so good?

OpenStudy (anonymous):

So when you write log it is already known that it's base is 10?

OpenStudy (solomonzelman):

When not indicated otherwise.

OpenStudy (anonymous):

Alright

OpenStudy (solomonzelman):

\[Log_24\]good with this one?

OpenStudy (solomonzelman):

\[\frac{Log_{10}4}{Log_{10}2}=Log_24~~~~~~~~~~~~~Good?\]

OpenStudy (anonymous):

Yes,but how about this side?\[\frac{ 2x }{\left( 2x-2 \right)}\]

OpenStudy (solomonzelman):

So far we have \[\frac{2x}{2(x-2)}=\log_24\]OK?

OpenStudy (anonymous):

Okay,then what's next?

OpenStudy (solomonzelman):

\[Log_bb=1~~~~~~~~~~identity/law\]

OpenStudy (solomonzelman):

\[Log(x \times y)=Log~x + Log~y\]\[Log(4)=Log(2 \times 2)=Log~2 + Log~2\]no matter what the mase is as long as it is the same base.

OpenStudy (anonymous):

If I am not mistaken,the above is the product law?

OpenStudy (solomonzelman):

Yes it is, you are correct, so far so good?

OpenStudy (anonymous):

Yup

OpenStudy (solomonzelman):

\[With~~base~~2:~~~~~~~~Log_2(4)=Log_2(2 \times 2)=Log_22+Log_22\]\[knowing~~that:~~~~\log_bb=1\]\[Log_22+Log_22=?\]you tell me

OpenStudy (anonymous):

Is it 1?

OpenStudy (solomonzelman):

\[Log_bb=1~~~~~~~~~Log_22=1\]\[Log_22+Log_22=1+1=2\]

OpenStudy (solomonzelman):

See?

OpenStudy (solomonzelman):

Do you see how i got \[Log_22+Log_22\]

OpenStudy (anonymous):

I get it now,thanks a lot

OpenStudy (isaiah.feynman):